OpenStudy (anonymous):

How do I integrate (4x^6+2x)^3(12x^5+1)dx by substitution?

OpenStudy (anonymous):

\[\int\limits_{-1}^{0}(4x^6+2x)^3(12x^5+1)\]

OpenStudy (anonymous):

\[=2\int\limits_{-1}^{0}\left( 2x^6+x \right)^3\left( 12x^5+1 \right)dx\] put \[2x^6+x=t,\left( 12x^5+1 \right)dx=dt\] find t when x=-1,and x=0 and integrate

OpenStudy (phi):

surj the 2 out front should be 2^3

OpenStudy (anonymous):

correct thanks

OpenStudy (anonymous):

So t = 2x^7/7 + x^2/2?

OpenStudy (anonymous):

I'm stuck...

OpenStudy (anonymous):

@surjithayer

OpenStudy (mathmale):

How are you stuck? Please share what you already know and what you need to know. "Stuck" is too general a term.

OpenStudy (anonymous):

Why can't I take a 2x out of the first set, why only a 2?

OpenStudy (anonymous):

whenx=-1 \[t=2(-1)^6+(-1)=2-1=1\] when x=0 \[t=2*0^6+0=0\] \[I=2^3\int\limits_{1}^{0}t^6dt=?\]